OpenStudy (anonymous):

3x less than or equal to 5x + 8

OpenStudy (anonymous):

\[3x \le 5x+8\] \[-2x \le 8\] \[x \ge -4\] (The sign switches from "less than or equal to" to "greater than or equal to" because we divided by a negative number.

OpenStudy (anonymous):

thanks